OpenStudy (anonymous):

a farmer divided his property among his three children as follows; the ratio of Sarah`s share to john`s share to Mary`s share is 2:3:5. if his property totals 75 acres , how many does each child receive?

OpenStudy (anonymous):

\[\frac{ 2 }{ 2+3+5 } = \frac{ ? }{ 75 }, where ? IsTheSarah'sShare\]

OpenStudy (anonymous):